OpenStudy (anonymous):

The current i in a charging circuit is given by i = 10e^(-100t) amperes. Find the charge due to the current after 10 milliseconds, if there is initially no charge in the circuit. So I did this: Integration: -[e^(-100t)] / 10 And when I subbed in t = 0.01, I got 0.0036 coulombs. However, the answer is 0.0632 coulombs. Is there any mistake with my integration, or my whole working in itself?

OpenStudy (anonymous):

Also, i = dq/dt, where q is the charge in the circuit, if that helps.

OpenStudy (anonymous):

If there's no problem with my integration, then why is it that I can't get the answer after subbing it in? :(

ganeshie8 (ganeshie8):

i think you're not subtracting the lower bound

ganeshie8 (ganeshie8):

check it once

OpenStudy (anonymous):

What do you mean by 'lower bound"?

ganeshie8 (ganeshie8):

\(\large Q = \int_0^{0.01} 10e^{-100t} dt \)

ganeshie8 (ganeshie8):

\(\large Q = \frac{-1}{10} e^{-100t} \Big|_0^{0.01} \)

ganeshie8 (ganeshie8):

\(\large Q = \frac{-1}{10} [e^{-1} - e^0] \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

Oh... I get it. But this question was given in the Indefinite Integral chapter(which is before the Definite Integral chapter), so is there a way to solve it without the bounds?

ganeshie8 (ganeshie8):

yeah sure :) it becomes an IVP

ganeshie8 (ganeshie8):

IVP = Initial Value Problem

ganeshie8 (ganeshie8):

\(\large Q(t) = \int 10e^{-100t} dt \) \(\large Q(t) = \frac{-1}{10} e^{-100t} + c \)

ganeshie8 (ganeshie8):

find the constant by plugging in the point (0, 0)

OpenStudy (anonymous):

Meaning, when t = 0?

ganeshie8 (ganeshie8):

yes, when t = 0, Q = 0 using that Initial Values, find the integration constant \(c\)

ganeshie8 (ganeshie8):

we got that infor from the question : ... if there is initially no charge in the circuit.

OpenStudy (anonymous):

So C = 0.1?

ganeshie8 (ganeshie8):

yup

ganeshie8 (ganeshie8):

so ur charge equation becomes : \(\large Q(t) = \frac{-1}{10} e^{-100t} + 0.1 \)

ganeshie8 (ganeshie8):

plugin t = 10ms, u should get the same answer..

OpenStudy (anonymous):

Oh. Lol. Okay, got it! Thanks a lot!

ganeshie8 (ganeshie8):

np.. u wlc :)